We heard that there was a 90% confidence interval for the population mean from 22.5 to 2am 8.3, determine the point estimate of the population and the margin of error associated with the information.


1
Expert's answer
2021-11-08T02:54:46-0500
"Confidence\ interval\ width"

"=2\times margin\ of\ error"

"margin\ of\ error=\dfrac{28.3-22.5}{2}=2.9"

"\bar{x}=22.5+2.9=25.4"

The sample mean "\bar{x}=25.4" is a point estimate of the population mean.

The margin of error is "2.9."